# Priors in Time: Missing Inductive Biases for Language Model Interpretability

### v1

@@ 1 Introduction
Understanding how representations in neural networks are updated over time is a key challenge for artificial intelligence and neuroscience (Rosenblatt, 1958; Elman, 1990). Large language models (LLMs) build intermediate representations incrementally over sequences of tokens. This provides a rich setting to study the temporal dynamics of representations. One class of interpretability tools, featurizers, aim to discretize these representations in a manner that is tractable for analysis and intervention (Mueller et al., 2025; Geiger et al., 2025). These tools serve a critical role in engineering applications—e.g., steering (Turner et al., 2024) and alignment (Lehalleur et al., 2025)—but also hold promise for testing hypotheses about how biological brains may implement computations underlying complex behaviors (Hosseini and Fedorenko, 2023; Tuckute et al., 2024a; Klindt et al., 2025).
Featurizers are built on hypothesized computational models of how concepts are encoded in neural network representations. Leading examples of computational models include the linear representation hypothesis (LRH; Elhage et al., 2022; Arora et al., 2018; Merullo et al., 2025; Hernandez et al., 2023), and corresponding tools include sparse autoencoders (SAEs; Gao et al., 2024; Cunningham et al., 2023; Bussmann et al., 2024; Rajamanoharan et al., 2024).
Critically, however, these models and methods make strong prior assumptions about the structure of concepts in LLMs. At the core of this paper is the following principle: interpretability methods should be driven by the behavior one is trying to explain. A mismatch between the structure of the data distribution and structural priors of interpretability tools can yield misleading or pathological results (Chanin et al., 2025; Hindupur et al., 2025; Costa et al., 2025; Bussmann et al., 2025; Bricken et al., 2023). For instance, mismatches can arise from using sparse coding (Olshausen and Field, 1997) to interpret language data. The sparsity penalty in sparse coding arises from the assumption that independent components combine to create the data; this was originally intended for images (Olshausen and Field, 1997). SAEs inherit the assumptions underlying sparse coding, so applying them to sequential data like language model activations may be pathological: language model activations are not independently distributed with respect to time.
In this work, we first study the pitfalls of applying SAEs to language model (LM) activations. In Fig.1, we use an illustrative example of Harry Potter to highlight the temporal structure of language and of language model activations. Language has temporal structure at multiple scales (Marslen-Wilson and Tyler, 1980; Thompson, 1999). At the sentence level, dependencies link words across time (Gibson et al., 2000; McElree et al., 2003), and upcoming words can be anticipated from local context, as formalized in expectation-based (probabilistic) accounts of language processing (Hale, 2001; Levy, 2008). Discourse and narratives impose structure over longer timescales, as in event boundary segmentation (Zacks et al., 2007; Baldassano et al., 2018). This structure is reflected in Fig. 1 as large numbers of shared attributes between nearby time points (i.e., local context), with only a few new attributes not present in local context added at a given time.
The temporal structure of language causes the resulting structure of LM activations, including strong local correlations over time (explored in Sec. 3). This structure raises the following questions: What are the temporal assumptions of existing SAEs, and do they capture the structure of LM activations? If there is a mismatch between the true and assumed data structure, as illustrated in Figs. 1 and 2, would existing SAEs be unable to reveal latent temporal structures? To overcome such limitations, how can we incorporate inductive biases truer to the temporal dynamics of natural language data into our featurizers? Motivated by these questions, this work makes the following contributions.
@@
  • Precise characterization of prior assumptions of existing SAEs. Adopting a Bayesian perspective on the SAE objective, we show that SAEs implicitly assume an independent and identically distributed (i.i.d.) prior across time on their latents, and thereby on concepts. As a consequence, SAEs fail to capture local correlations in LM activations.
@@
  • Highlighting temporal structure in language model (LM) representations. We demonstrate rich temporal dynamics in LM representations, which show nonstationarity through systematic growth in conceptual dimensions with time, time-varying activation correlations, and strong correlations with recent context, that are at odds with the implicit priors of SAEs (Fig. 1).
@@
  • Incorporating data-informed temporal inductive biases into a novel SAE architecture. Using empirically observed temporal structure in LM activations, we design a novel SAE architecture–called Temporal Feature Analysis–which decompose the data into two components: a predictive component, which captures slow-moving contextual information, and a novel component, which captures fast-moving stimulus-driven information.
@@
  • Using Temporal Feature Analysis to reveal temporal structures in LM representations. We demonstrate Temporal Feature Analysis’s ability to extract temporal structures by studying linguistic input with temporal structure at different scales: garden path sentences, stories, and in-context dialogue. Specifically, we show the extracted predictive codes: (1) correctly parse garden path sentences, (2) decompose stories into events, (3) capture precise structure of in-context representations and allow counterfactual interventions, and (4) delineate user-assistant interactions through smooth trajectories. Meanwhile, the novel codes allow automatic interpretability pipelines, akin to SAEs.
@@ 2 Preliminaries
Sparse dictionary learning (Olshausen and Field, 1996; 1997) expresses data as a sparse linear combination of dictionary elements, where both the weights and dictionary are learned from data. Intuitively, the dictionary behaves as a data-adaptive overcomplete basis; i.e., it typically has more elements that the dimension of ambient space. The optimization problem involved in this framework is $\operatorname*{arg\,min}_{\bm{D},\bm{z}}\frac{1}{N}\sum_{i=1}^{N}\|\bm{x}_{i}-\bm{D}\bm{z}_{i}\|_{2}^{2}+\lambda\mathcal{R}(\bm{z}_{i})$, where $\mathcal{R}(\cdot)$, typically chosen to be the $\ell_{1}$-norm, is a sparsity-inducing regularizer. Sparsity assists in picking the fewest most relevant dictionary atoms to explain a given data point.
@@
SAEs aim to disentangle (Bengio et al., 2013; Higgins et al., 2018; Olah, 2023) neural network activations into human-interpretable concepts (Cunningham et al., 2023; Bricken et al., 2023). Specifically, SAEs transform their inputs (i.e., neural network activations) into a latent representation which is encouraged to be sparse. As shown by Hindupur et al. (2025), this is achieved by solving the sparse coding problem using a specific parametric form for the sparse codes:
@@ 3 Temporal Structure in Language Model Activations
Figure 2: Temporal structure of LLM activations reveals nonstationarity. We use Pile samples (Monology, 2021) to analyze temporal structure from activations of two pretrained LMs, comparing it to a surrogate signal that is stationary in nature (see App. A.1.4). (a, e) Intrinsic dimension of model activations and stationary surrogate. (b, f) Autocorrelations $A(\bm{x}_{t},\bm{x}_{t-\tau})$ as a function of sequence position ($t$) and lag ($\tau$). (c, g) Autocorrelation of the stationary surrogate. (d, h) Variance explained by projecting current representation $\bm{x}_{t}$ onto past context window $\{\bm{x}_{t-1},\dots,\bm{x}_{t-w}\}$ with different sizes $w$, along with a baseline. Results consistently show representations getting ‘denser’ over time and being significantly more structured than a stationary surrogate.
To contextualize the prior assumptions made by SAEs about temporal structure in LMs’ activations, we first perform an empirical characterization of such temporal structure in pretrained LMs. Specifically, since LMs are trained to generate coherent text by learning the distribution of natural language, one can expect their representations capture the rich phenomenology of its sequential structure (Elman, 1990); indeed, recent work has in fact found LM representations to be predictive of human neural recordings during language comprehension (Hosseini et al., 2024; Schrimpf et al., 2021; Tuckute et al., 2024b; Hong et al., 2024; Georgiou et al., 2023). Motivated by this, we perform two experiments relevant to our discussion (see Fig. 2): (i) measuring intrinsic dimensionality—an approximation of the number of concepts necessary to explain the data, which can be expected to increase in a monotonic manner with time (Zhong et al., 2024; Can, 2025; Barak and Tsodyks, 2014; Meister et al., 2021)—and (ii) signal nonstationarity—which assesses whether model activations reflect the contextual relations between phrases of a passage  (Zacks et al., 2007).
@@
Subplots (b), (f) show the autocorrelation of model activations (App. Sec. A.1.1), which is noticeably different at different sequence positions (x-axis), while the stationary surrogate, as expected, shows nearly position-invariant autocorrelation values (subplots (c), (g)). This finding is a clear signature of time-dependent correlation structure, and therefore of nonstationarity. To evaluate this further, in subplots (d), (h), we project representations of token $\bm{x}_{t}$ at a given time $t$ onto the PCA basis obtained from a context window $C(\bm{x},t,w)=[\bm{x}_{t-w},\ldots,\bm{x}_{t-1}]$ with different sizes $w$. These subplots show that up to $\sim$$80\%$ variance of $\bm{x}_{t}$ is explained by $C(\bm{x},t,w)$ with $w$ up to 100; this highlights strong cross-temporal correlations. This observation indicates that significant variance in the representation at time $t$, $\bm{x}_{t}$, can be predicted (expressed) using representations from the past context $\bm{x}_{<t}$.
@@ 4 Temporal Priors of Sparse Autoencoders
We now state the prior assumptions made by existing SAEs regards sequential structure in an input, contrasting these assumptions with the empirical results shown in Sec. 3. Specifically, building on the arguments used by Olshausen and Field (1997) to formalize the problem of sparse coding, we note that the SAE training objective (Eq. 1) can be interpreted from a Bayesian lens: minimize the negative log posterior $\operatorname*{arg\,min}_{\{\bm{z}_{t}\}}-\log P(\bm{z}_{1},\dots,\bm{z}_{T}\mid\bm{x}_{1},\dots,\bm{x}_{T})$ of the data, which, by Bayes’ rule, can be written as the sum of log likelihood (MSE) and log prior (the regularizer $\mathcal{R}$). From this lens, SAEs’ prior assumptions on sequential structure in LM activations can be described as follows.
@@
SAEs thus assume that sparsity—which, in their underlying generative model of activations corresponds to the number of concepts necessary for explaining the data (Bricken et al., 2023; Elhage et al., 2022)—remains approximately constant over time. This again does not align with the increasing dimensionality of representations observed in LM activations (see Fig. 2a,e). Correspondingly, if enough concepts aggregate over context such that a model’s activations become ‘denser’ than the assumed sparsity budget, the assumption of time-invariant sparsity implies SAEs can fail to capture the temporal structure inherent in language, as formally stated below.
@@
Intuitively, the claim above says that even if an SAE achieves low reconstruction error, its specific geometric assumptions—use of independently firing directions and restricted intrinsic dimensionality of data—will lead to a “splitting” of the input data into several regions such that different sets of latents will activate for even locally nearby points. Consequently, if the sparsity budget is too little or if model activations are more dense than the assumed budget, then local geometry present in the activations will be lost in the SAE latent codes—as is arguably already observed empirically with phenomena like feature splitting (Chanin et al., 2025; Bussmann et al., 2025; Bricken et al., 2023). For the context of our paper, as per our results in Fig. 2 (a,e), temporal correlations can increase activations’ intrinsic dimensionality and, when viewed from the lens of the claim above, we arrive at the conclusion that SAEs can fail to capture temporal structure present in activations. We empirically validate this claim in Sec. 6. More broadly, we note there is growing evidence that neural network representations possess rich geometric structures across a broad set of modalities (Gurnee et al., 2025; Fel et al., 2025; Park et al., 2025; Engels et al., 2024; Costa et al., 2025; Kantamneni and Tegmark, 2025; Pearce et al., 2025). As per Prop. 4.2, the use of SAEs for interpreting behaviors that invoke such structures may yield misleading explanations and should be approached carefully.
@@ 5 Temporal Feature Analysis: Towards Predictive Approaches for Interpreting Neural Representations
Figure 4: Schematic of Temporal Feature Analysis. Temporal Feature Analyzers decompose activations $\bm{x}_{t}$ into two components: a predictable component, obtained by projecting $\bm{x}_{t}$ onto a context direction (derived from the past $\bm{x}_{<t}$ using attention), and a novel component orthogonal to the predictable component, capturing new information emerging at time $t$ (the novel code $\bm{z}_{n,t}$ is sparse).
As stated in Sec. 2, sparse coding, a framework designed in computational neuroscience to understand neural representations in biological brains (Olshausen and Field, 1996; 1997), inspired SAEs as a framework for interpreting artificial neural networks (Bricken et al., 2023; Cunningham et al., 2023). In fact, the parallels between these communities can be made deeper: motivated by observations of intricate geometry of neural representations derived out of multi-dimensional population analyses (Khona and Fiete, 2022; Nogueira et al., 2023; Sohn et al., 2019), there were calls in computational neuroscience to discard the limiting reduction assumed in sparse coding that computations occur via a set of independently firing, monosemantic features (Eichenbaum, 2018; Saxena and Cunningham, 2019; Barack and Krakauer, 2021; Seung, 1996; Chung and Abbott, 2021)—similar to our arguments in Sec. 34 (and results that follow in Sec. 6). Correspondingly, a need for more structured protocols was suggested (Eichenbaum, 2018; Barack and Krakauer, 2021; Chen, 2019), leading to methods that were motivated by the generative process of the behavior one is trying to explain (Schneider et al., 2023; Chen et al., 2018; Berkes and Wiskott, 2005; Wiskott and Sejnowski, 2002; Linderman et al., 2017; Yu et al., 2008). We argue a similar paradigm shift is needed in language model interpretability: given that we train models to learn the distribution of highly structured data, we ought to embrace the fact that neural network activations can exhibit intricate geometrical organization. These geometrical objects may in fact be the units of computation that are necessary for describing model computation (Perich et al., 2025; Eichenbaum, 2018; Modell et al., 2025), and hence we must design interpretability approaches to identify them. In what follows, as an attempt to qualify our arguments, we propose one such approach, titled Temporal Feature Analysis, that focuses on the temporal structure of LM activations.
@@
In computational neuroscience, when analyzing data from dynamical domains (e.g., audio, language, or video), a commonly made assumption is that there is contextual information present in the recent history that informs the next state—this part of the signal is deemed predictable (Chen et al., 2025; Millidge et al., 2024), slow-changing (Berkes and Wiskott, 2005), invariant (Olshausen and Cadieu, 2007), or dense (Tasissa et al., 2022). Meanwhile, the remaining signal corresponds to new bits of information added by the observed state at the next timestep—this part can be deemed novel, fast-changing, variant, or sparse with respect to the context. We argue LM activations are amenable to a similar generative model. Specifically, our observations in Sec. 3 show that activations $\bm{x}_{t}$ at time $t$ are strongly correlated with the context and can be decomposed into two such parts as well. Motivated by this, we propose the following model of LM activations:
@@
where $\bm{x}_{p,t}=\bm{D}\bm{z}_{p,t}$ denotes a predictable component of the signal that captures the correlations of $\bm{x}_{t}$ with past data $\left\{\bm{x}_{<t}\right\}$, while $\bm{x}_{n,t}=\bm{D}\bm{z}_{n,t}$ denotes a novel component that represents new information added by the current token $\bm{x}_{t}$. To obtain $\bm{z}_{p,t}$, we project $\bm{x}_{t}$ onto $\{\bm{x}_{<t}\}$ to explain the predictable variance in $\bm{x}_{t}$ as a convex combination of past data. Specifically, we use a self-attention layer $f$ on top of a single ReLU layer, yielding $\bm{z}_{p,t}=f(\ \{\bm{x}_{1},\dots,\bm{x}_{t-1}\},\bm{x}_{t})$. Meanwhile, $\bm{z}_{n,t}=\tilde{f}(\bm{x}_{t},\bm{z}_{p,t})=\sigma(\bm{D}^{T}(\bm{x}_{t}-\bm{D}\bm{z}_{p,t}))$ captures the residual component of the code, which is not correlated with the past (note that $\sigma$ is the nonlinearity). We use a standard SAE encoder (either TopK or BatchTopK) to instantiate $\sigma$, applying it to $\bm{x}_{t}-\bm{D}\bm{z}_{p,t}$ to derive $\bm{z}_{n,t}$. See Fig. 4 for an overall schematic of the encoding process. The overall learning objective in Temporal Feature Analysis is as follows:

### v2

@@ 1 Introduction
Given the success of Language Models (LMs) (bubeck2023sparks; geminigold), there is growing interest in understanding how such models incrementally update over sequences of tokens to exhibit complex behaviors (murthy2025inside; lindsey2025emergent; lindsey2025biology; lepori2025just; bigelow2025belief; tuckute2024driving; klindt2025superposition). Interpretability research aims to make such analyses tractable, offering tools for hypothesis design, testing, and intervention based on evaluation of intermediate activations (geiger2025abstraction; sharkey2025open; bereska2024mechanistic). Often, such work builds on hypothesized computational models of how concepts are encoded in a neural network’s representations, e.g., the linear representation hypothesis (LRH) (elhage2022superposition; arora2018linear), correspondingly motivating tools such as sparse autoencoders (SAEs) (gao2024scaling; cunningham2023sparse) for unsupervised extraction of a dictionary of vectors that (ideally) mediate human-interpretable concepts (mueller2025questrightmediatorsurveying).
A central challenge in “bottom-up” approaches to interpretability, like SAEs, is the mismatch between the assumptions of their underlying implementational account and the precise behavior or computation they intend to explain (jonas2017could; geiger2025abstraction; costa2025flat) (see Fig. 1). For instance, since LRH posits that different concepts correspond to directions in activation space that can be independently manipulated, it implicitly claims the data distribution can be factorized into independently varying latent variables (allen2024unpicking). This mismatch between the structure of the data distribution and strong priors codified in LRH can lead to misleading or pathological explanations when using SAEs to understand neural networks (chanin2025absorptionstudyingfeaturesplitting; bricken2023monosemanticity; hindupur2025projecting). This raises a set of critical questions for using SAEs to interpret models trained on sequential data like language. Specifically, since language exhibits rich temporal structure at multiple scales (marsenwilson-1980; thompson1999temporal)—e.g., sentences contain dependencies that link words across time (gibson2000dependency; mcelree2003memory), upcoming words can be anticipated from context (hale2001probabilistic; levy2008expectation), and discourse imposes structure over longer timescales through phenomena like event boundaries (zacks2007event; baldassano2017discovering)—one can ask what assumptions about temporal structures do SAEs make? How do these assumptions align with the actual temporal structure present in a LM’s activations?
@@
This work.
Building on the Bayesian interpretation of sparse coding (olshausen1996emergence; olshausen1997sparse)—the framework that motivates SAEs—we rephrase the optimization objective of SAEs as a MAP (maximum a posteriori) estimation problem. This allows us to make explicit prior assumptions about temporal structure embedded in SAEs, showing they implicitly assume concepts are uncorrelated across time and the number of concepts necessary to explain an activation is time-invariant—that is, the information present at each token position is independent of information at other positions and uniformly distributed. As we empirically show, these assumptions stand in stark contrast to the actual temporal structure present in language and language model representations, and can result in empirically observed pathologies in SAEs, such as feature splitting (bricken2023monosemanticity; chanin2025absorptionstudyingfeaturesplitting; bussmann2025learning). These results then motivate us to draw a broader parallel between SAEs and computational neuroscience approaches for understanding neural data. Specifically, population-level analyses of neural recordings have revealed that representations often lie on structured manifolds (khona2022attractor; nogueira2023geometry; sohn2019bayesian), challenging the reductionist assumption in sparse coding that computations occur via independently firing, monosemantic features (eichenbaum2018barlow; saxena2019towards; barack2021two). This motivated a paradigm shift towards more structured analysis protocols—methods designed around the generative process of the behavior one aims to explain (schneider2023learnable; chen2018sparse). Motivated by this and similar findings of intricate geometrical structure in neural network representations (fel2025into; gurnee2025when; modell2025origins), we propose Temporal Feature Analysis, a new protocol for interpreting language model activations that incorporates explicit inductive biases about temporal structure. Our approach decomposes activations at each timestep into two orthogonal components: a predictable component, obtained by projecting current representations onto past context using a learned attention mechanism, and a novel component, representing residual information orthogonal to the predictable component. That is, we assume the novel component—not the total representation—is uncorrelated over time. This allows correlations between total codes and hence enables our method to capture the temporal structure of LM activations. Overall, we make the following contributions in this paper.
@@
  • Highlighting temporal structure in language model (LM) representations (Sec. 3). We demonstrate rich temporal dynamics in LM representations, which show nonstationarity through systematic growth in conceptual dimensions with time, time-varying activation correlations, and strong correlations with recent context, that are at odds with the implicit priors of SAEs.
  • Precise characterization of prior assumptions of existing SAEs (Sec. 4). Adopting a Bayesian perspective on the SAE objective, we show that SAEs implicitly assume an independent and identically distributed (i.i.d.) prior across time on their latents, and thereby on concepts. As a consequence, SAEs fail to capture local correlations in LM activations.
@@
  • Incorporating data-informed temporal inductive biases into a novel SAE architecture (Sec. 5). Using empirically observed temporal structure in LM activations, we design a novel interpretability protocol–called Temporal Feature Analysis–which decomposes the model activation into two components: a predictive component, which captures slow-moving contextual information, and a novel component, which captures fast-moving stimulus-driven information.
@@
  • Using Temporal Feature Analysis to reveal temporal structures in LM representations (Sec. 6). We show Temporal Feature Analysis’s ability to extract temporal structures by studying linguistic input with temporal structure at different scales: garden path sentences, stories, and in-context dialogue. Specifically, we show the extracted predictive codes: (1) correctly parse garden path sentences, (2) decompose stories into events, (3) capture precise structure of in-context representations, and (4) delineate user-assistant interactions through smooth trajectories. Meanwhile, the novel codes allow automatic interpretability pipelines, akin to SAEs.
@@ 2 Preliminaries
Sparse dictionary learning (olshausen1996emergence; olshausen1997sparse) expresses data as a sparse linear combination of dictionary elements, where both the weights and dictionary are learned from data. Intuitively, the dictionary behaves as a data-adaptive overcomplete basis; i.e., it typically has more elements that the dimension of ambient space. The optimization problem involved in this framework is $\operatorname*{arg\,min}_{\bm{D},\bm{z}}\frac{1}{N}\sum_{i=1}^{N}\|\bm{x}_{i}-\bm{D}\bm{z}_{i}\|_{2}^{2}+\lambda\mathcal{R}(\bm{z}_{i})$, where $\mathcal{R}(\cdot)$, typically chosen to be the $\ell_{1}$-norm, is a sparsity-inducing regularizer. Sparsity assists in picking the fewest most relevant dictionary atoms to explain a given data point.
@@
SAEs aim to disentangle (bengio2013representation; higgins2018towards; olahreps) neural network activations into human-interpretable concepts (cunningham2023sparse; bricken2023monosemanticity). Specifically, SAEs transform their inputs (i.e., neural network activations) into a latent representation which is encouraged to be sparse. As shown by hindupur2025projecting, this is achieved by solving the sparse coding problem using a specific parametric form for the sparse codes:
@@ 3 Temporal Structure in Language Model Activations
Figure 2: Temporal structure of LLM activations reveals nonstationarity. We use Pile samples (monology2021pile-uncopyrighted) to analyze temporal structure from activations of two pretrained LMs, comparing it to a surrogate signal that is stationary in nature (see App. A.1.4). (a, e) Intrinsic dimension of model activations and stationary surrogate. (b, f) Autocorrelations $A(\bm{x}_{t},\bm{x}_{t-\tau})$ as a function of sequence position ($t$) and lag ($\tau$). (c, g) Autocorrelation of the stationary surrogate. (d, h) Variance explained by projecting current representation $\bm{x}_{t}$ onto past context window $\{\bm{x}_{t-1},\dots,\bm{x}_{t-w}\}$ with different sizes $w$, along with a baseline. Results consistently show representations getting ‘denser’ over time and being significantly more structured than a stationary surrogate.
To contextualize the prior assumptions made by SAEs about temporal structure in LMs’ activations, we first perform an empirical characterization of such temporal structure in pretrained LMs. Specifically, since LMs are trained to generate coherent text by learning the distribution of natural language, one can expect their representations capture the rich phenomenology of its sequential structure (elman1990finding); indeed, recent work has in fact found LM representations to be predictive of human neural recordings during language comprehension (hosseini2024universality; schrimpf2021neural; tuckute2024driving; hong2024scale; georgiou2023using). Motivated by this, we perform two experiments relevant to our discussion (see Fig. 2): (i) measuring intrinsic dimensionality—an approximation of the number of concepts necessary to explain the data, which can be expected to increase in a monotonic manner with time (zhong2024random; can2025statistical; barak2014working; meister2021revisiting)—and (ii) signal nonstationarity—which assesses whether model activations reflect the contextual relations between phrases of a passage  (zacks2007event).
@@
Subplots (b), (f) show the autocorrelation of model activations (App. Sec. A.1.1), which is noticeably different at different sequence positions (x-axis), while the stationary surrogate, as expected, shows nearly position-invariant autocorrelation values (subplots (c), (g)). This finding is a clear signature of time-dependent correlation structure, and therefore of non-stationarity. We quantify the similarity of a representation with its context in subplots (d), (h). Specifically, we project representations of token $\bm{x}_{t}$ at a given time $t$ onto the subspace spanned by preceeding representations in the context $\{\bm{x}_{<t}\}$. These subplots show that up to $80\%$ variance of $\bm{x}_{t}$ is explained by a context of 500 tokens, further highlighting strong cross-temporal correlations. Significant variance in the representation at time $t$, $\bm{x}_{t}$, can be predicted (expressed) using representations from the past context.
@@ 4 Temporal Priors of Sparse Autoencoders
We now state the prior assumptions made by existing SAEs regards sequential structure in an input, contrasting these assumptions with the empirical results shown in Sec. 3. Specifically, building on the arguments used by olshausen1997sparse to formalize the problem of sparse coding, we note that the SAE training objective (Eq. 1) can be interpreted from a Bayesian lens: minimize the negative log posterior $\operatorname*{arg\,min}_{\{\bm{z}_{t}\}}-\log P(\bm{z}_{1},\dots,\bm{z}_{T}\mid\bm{x}_{1},\dots,\bm{x}_{T})$ of the data, which, by Bayes’ rule, can be written as the sum of log likelihood (MSE) and log prior (the regularizer $\mathcal{R}$). From this lens, SAEs’ prior assumptions on sequential structure in LM activations can be described as follows.
@@
SAEs thus assume that sparsity—which, in their underlying generative model of activations corresponds to the number of concepts necessary for explaining the data (bricken2023monosemanticity; elhage2022superposition)—remains approximately constant over time. This again does not align with the increasing dimensionality of representations observed in LM activations (see Fig. 2a,e). Correspondingly, if enough concepts aggregate over context such that a model’s activations become ‘denser’ than the assumed sparsity budget, the assumption of time-invariant sparsity implies SAEs can fail to capture the temporal structure inherent in language, as formally stated below.
@@
Intuitively, the claim above says that even if an SAE achieves low reconstruction error, its specific geometric assumptions—use of independently firing directions and restricted intrinsic dimensionality of data—will lead to a “splitting” of the input data into several regions such that different sets of latents will activate for even locally nearby points. Consequently, if the sparsity budget is too little or if model activations are more dense than the assumed budget, then local geometry present in the activations will be lost in the SAE latent codes—as is arguably already observed empirically with phenomena like feature splitting (chanin2025absorptionstudyingfeaturesplitting; bussmann2025learning; bricken2023monosemanticity). For the context of our paper, as per our results in Fig. 2 (a,e), temporal correlations can increase activations’ intrinsic dimensionality and, when viewed from the lens of the claim above, we arrive at the conclusion that SAEs can fail to capture temporal structure present in activations. We empirically validate this claim in Sec. 6. More broadly, we note there is growing evidence that neural network representations possess rich geometric structures across a broad set of modalities (gurnee2025when; fel2025into; park2025iclrincontextlearningrepresentations; engels2024not; costa2025flat; kantamneni2025language; pearce2025tree). As per Prop. 4.2, the use of SAEs for interpreting behaviors that invoke such structures may yield misleading explanations and should be approached carefully.
@@ 5 Temporal Feature Analysis: Towards Predictive Approaches for Interpreting Neural Representations
Figure 4: Schematic of Temporal Feature Analysis. Temporal Feature Analyzers decompose activations $\bm{x}_{t}$ into two components: a predictable component, obtained by projecting $\bm{x}_{t}$ onto a context direction (derived from the past $\bm{x}_{<t}$ using attention), and a sparse, novel component orthogonal to the predictable component that captures new information seen at time $t$.
As stated in Sec. 2, sparse coding, a framework designed in computational neuroscience to understand neural representations in biological brains (olshausen1996emergence; olshausen1997sparse), inspired SAEs as a framework for interpreting artificial neural networks (bricken2023monosemanticity; cunningham2023sparse). In fact, the parallels between these communities can be made deeper: motivated by observations of intricate geometry of neural representations derived out of multi-dimensional population analyses (khona2022attractor; nogueira2023geometry; sohn2019bayesian), there were calls in computational neuroscience to discard the limiting reduction assumed in sparse coding that computations occur via a set of independently firing, monosemantic features (eichenbaum2018barlow; saxena2019towards; barack2021two; seung1996brain; chung2021neural)—similar to our arguments in Sec. 34 (and results that follow in Sec. 6). Correspondingly, a need for more structured protocols was suggested (eichenbaum2018barlow; barack2021two; chen2019sparse), leading to methods that were motivated by the generative process of the behavior one is trying to explain (schneider2023learnable; chen2018sparse; berkes2005slow; wiskott2002slow; linderman2017bayesian; yu2008gaussian). We argue a similar paradigm shift is needed in language model interpretability: given that we train models to learn the distribution of highly structured data, we ought to embrace the fact that neural network activations can exhibit intricate geometrical organization. These geometrical objects may in fact be the units of computation that are necessary for describing model computation (perich2025neural; eichenbaum2018barlow; modell2025origins), and hence we must design interpretability approaches to identify them. In what follows, as an attempt to qualify our arguments, we propose one such approach, titled Temporal Feature Analysis, that focuses on the temporal structure of LM activations.
@@
In computational neuroscience, when analyzing data from dynamical domains (e.g., audio, language, or video), a commonly made assumption is that there is contextual information present in the recent history that informs the next state—this part of the signal is deemed predictable (chen2025context; millidge2024predictive), slow-changing (berkes2005slow), invariant (olshausen2007learning), or dense (tasissa2022improvingdiscriminativereconstructionsimultaneous). Meanwhile, the remaining signal corresponds to new bits of information added by the observed state at the next timestep—this part can be deemed novel, fast-changing, variant, or sparse with respect to the context. We argue LM activations are amenable to a similar generative model. Specifically, our observations in Sec. 3 show that activations $\bm{x}_{t}$ at time $t$ are strongly correlated with the context and can be decomposed into two such parts as well. Motivated by this, we propose the following model of LM activations:
@@
where $\bm{x}_{p,t}=\bm{D}\bm{z}_{p,t}$ denotes a predictable component of the signal that captures the correlations of $\bm{x}_{t}$ with past data $\left\{\bm{x}_{<t}\right\}$, while $\bm{x}_{n,t}=\bm{D}\bm{z}_{n,t}$ denotes a novel component that represents new information added by the current token $\bm{x}_{t}$. To obtain $\bm{z}_{p,t}$, we project $\bm{x}_{t}$ onto $\{\bm{x}_{<t}\}$ to explain the predictable variance in $\bm{x}_{t}$ as a convex combination of past data. Specifically, we use a self-attention layer $f$ on top of a single ReLU layer, yielding $\bm{z}_{p,t}=f(\ \{\bm{x}_{1},\dots,\bm{x}_{t-1}\},\bm{x}_{t})$. Meanwhile, $\bm{z}_{n,t}=\tilde{f}(\bm{x}_{t},\bm{z}_{p,t})=\sigma(\bm{D}^{T}(\bm{x}_{t}-\bm{D}\bm{z}_{p,t}))$ captures the residual component of the code, which is not correlated with the past (note that $\sigma$ is the nonlinearity). We use a standard SAE encoder (either TopK or BatchTopK) to instantiate $\sigma$, applying it to $\bm{x}_{t}-\bm{D}\bm{z}_{p,t}$ to derive $\bm{z}_{n,t}$. See Fig. 4 for an overall schematic of the encoding process. The learning objective in Temporal Feature Analysis follows.